• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Ole Hansen

    Log entry time 16:04:21 on September 1,2009

    Entry number 287350

    keyword=Day Shift Summary

    8:00 Start shift with production HAPPEX data taking.
    8:50 PSS problem takes the Hall down at 8:50.
    9:00-12:00 PSS troubleshooting. Simultaneously, cryo work (bringing ESR compressor online), requiring target and magnet warmup.
    12:00-14:00 Target cooldown. Controlled access for target maintenance and minor other work.
    14:00-15:20 Restoring beam. Raster checks. Target motion. Discussions of run plan.
    15:20-16:00 Beam gone for Hall C beamline tuning.